# Less is More: Hop-Wise Graph Attention for Scalable and Generalizable Learning on Circuits

### 06 Mar 2024 v2

@@ Abstract
I Introduction
@@
Recent years have seen a surge of interest in machine learning (ML) for electronic design automation (EDA), which holds great potential in achieving faster design closure and minimizing the need for extensive human supervision [10]. In particular, graph neural networks (GNNs) have become increasingly popular in the EDA community due to their ability to encode graph-structured data such as gate-level netlists into compact representations, which can be used for a multitude of downstream EDA applications, including quality of results (QoR) prediction and functional reasoning [13, 18].
@@
However, scaling GNN training to large graphs is a notoriously challenging problem, which poses a serious concern on the practical benefit of GNNs on large-scale EDA problems. On the one hand, unlike common datasets on social networks and molecular graphs, which consist of either a few large graphs or a large number of small graphs, the circuit datasets may contain numerous graphs with each graph containing hundreds of thousands of nodes. For instance, the OpenABC-D benchmark provides $870$k gate-level netlists, where each netlist consists of up to $240$k logic gates [6]. Thus, training GNNs on such a large-scale circuit dataset is even more challenging than other graph-based applications. On the other hand, the mainstream GNN models are built upon the conventional message-passing paradigm among nodes, which learns node representations through a recursive node-wise aggregation scheme shown in Figure 1(b). As a consequence, it is nontrivial to perform efficient distributed GNN training due to the node dependencies in a graph structure.
@@
Apart from the scalability challenge, it is also underexplored how to make GNNs generalizable across different circuit designs. Although there are many customized GNNs previously proposed for various EDA applications, their model backbones mainly follow classic GNNs such as GCN [11] and GraphSAGE [9], which are not necessarily suitable for circuit problems. Consider a task of identifying functional blocks within circuits [18]. As distinct functional blocks may have different depths, the number of hops to be considered varies across nodes, which cannot be easily captured by common GNNs. Moreover, the high-order structures of functional blocks are also important yet ignored by the aforementioned GNN models. As a result, existing GNNs for EDA tasks often struggle to learn the intrinsic and critical information from complex circuit graphs, resulting in limited generalizability to unseen designs.
@@
Figure 1: Comparison of HOGA and conventional GNNs — (a) An example graph for illustration; (b) Computation graph of GNNs; (c) Computation graph of our proposed approach, HOGA.
@@
In literature, improving the scalability and enhancing the generalizability of GNNs on circuits are largely viewed as two orthogonal directions and seldomly explored together. While there is a rich body of prior studies focusing on scaling GNNs to large graphs with distributed training, they either incur massive communication overhead among GPUs, or rely on heuristic graph sampling algorithms that may lose critical structural information [14], which in turn degrades the model generalizability. In the context of learning generalizable GNNs for circuit problems, Wang et al. [17] leveraged graph contrastive learning to pretrain GNNs and adopted multiple neighbor aggregation functions to learn the functionalities of logic gates, which comes with additional computational costs and thus worsens the model scalability.
@@
Motivated by the limitations of conventional GNNs on circuit designs, we propose a novel hop-wise attention approach, named HOGA, to improve both the scalability and generalizability of circuit representation learning. As shown in Figure 1(c), HOGA adopts a hop-wise aggregation scheme, which precomputes the hop-wise features and only uses those features to learn node representations through a gated self-attention module. Since the final representation per node depends solely on its own hop-wise features, there are no dependencies between different nodes during training, making it easy to scale HOGA training in a distributed manner. Moreover, the gated self-attention module enables HOGA to adaptively aggregate important and high-order features from different hops per node, rendering it generalizable across different circuit designs.
@@
It is noteworthy that HOGA is a flexible approach for learning circuit representations, which can be integrated with other customization techniques previously proposed for various downstream EDA tasks. To demonstrate the viability and flexibility of HOGA, we consider two representative circuit problems: (1) QoR prediction – We focus on predicting the optimized gate count after logic synthesis on the OpenABC-D benchmark, which is generated from various circuit designs as well as synthesis recipes, and is one of the largest open-sourced circuit datasets; (2) Functional reasoning – We follow the most challenging setting in Gamora [18] by identifying functional blocks on gate-level netlists after technology mapping. Our experiments show that HOGA not only outperforms prior GNNs on unseen designs in both applications, but also scales comfortably to large-scale circuits. We summarize our main technical contributions as follows:
@@
$\bullet$ To the best of our knowledge, we are the first to introduce a scalable and generalizable model for circuit representation learning, which is achieved by a coarse-grained message-passing scheme based on hop-wise attention.
@@
$\bullet$ By precomputing hop-wise features, HOGA avoids recursive neighbor aggregation during training. This enables HOGA to learn each node representation independently and facilitates massive parallelization for distributed training. As a result, the training time of HOGA almost linearly decreases with an increase in computing resources.
@@
II Preliminary and Motivation
II-A Graph Neural Networks at Scale
@@
$m_{i}^{(l)}=\text{AGGREGATE}^{(l)}(\{x_{j}^{(l-1)}\;|\;j\in\mathcal{N}(i)\})$ (1)
$x_{i}^{(l)}=\text{COMBINE}^{(l)}(m_{i}^{(l)},x_{i}^{(l-1)})$ (2)
where $x_{i}^{(l)}$ denotes the node feature vector at the $l$-th layer, $\mathcal{N}(i)$ represents the set of $1$-hop neighbors of node $i$, AGGREGATE can be any permutation-invariant operator such as MEAN or MAX pooling, and the goal of COMBINE function is to update node features based on the aggregated features from neighbors. After stacking $L$ GNN layers, the output features $\{x_{i}^{(L)}\;|\;i\in\mathcal{V}\}$ serve as the node representations that are used for downstream EDA applications.
However, producing a single node representation $x_{i}^{(L)}$ involves a recursive neighbor aggregation scheme, which can be straightforwardly derived by substituting Equation (1) into Equation (2). More concretely, let $d$ be the average node degree, there are $O(d^{L})$ nodes required to obtain $x_{i}^{(L)}$, which increases exponentially with the number of layers and is known as the “neighbor explosion” issue. Hence, GNNs generally suffer from poor scalability to large graphs, prohibiting their applications in large-scale circuit problems.
@@
As a remedy, distributed training has been increasingly used for scaling GNNs to large graphs. However, handling the dependencies in a graph structure incurs significant communication overhead and work imbalance, hindering efficient distributed GNN training. To tackle this challenge, numerous efforts have been devoted to increase GNN training parallelism by adopting node-wise [9], layer-wise [4], and graph-wise sampling [20]. While these methods demonstrate promising results on common graph datasets such as social networks, we argue that they are unsuitable for circuit graphs, since the sampling algorithms may entirely break the design functionality and lead to poor accuracy. This is empirically confirmed in Section IV-C. In contrast, we introduce a novel hop-wise aggregation scheme, which independently learns each node representation based on its own hop-wise features. This renders our model embarrassingly parallel and greatly facilitates distributed training.
II-B Generalizable Graph Learning in EDA
@@
Since realistic circuit designs may have different sizes or originate from distinct domains, generalization capability is crucial for deploying graph learning models in EDA applications. To this end, prior work has developed many customization techniques, particularly for GNNs in learning circuit representations. For instance, Ustun et al. proposed a customized GraphSAGE model by distinguishing predecessors and successors in a graph, which substantially outperforms the vanilla GraphSAGE on learning operation mapping patterns [15]. Later, Zhang et al. [21] and Guo et al. [8] introduced customized GNNs tailored for power inference and timing prediction tasks respectively, via sequentially updating node representations. More recently, Wu et al. presented a multi-task graph learning framework for functional reasoning. While these methods showcase promising results on the respective tasks, their GNN backbones are built upon the conventional message-passing paradigm, which cannot capture critical high-order structures formed by multiple nodes. It is worth noting that high-order structures are crucial for many circuit problems, such as functional reasoning. Specifically, identifying a half adder and a full adder within circuits requires to learn different high-order structures that correspond to their respective functional blocks [18]. Therefore, the aforementioned methods still cannot capture the intrinsic circuit information, leading to either restrictions to specific tasks or limited generalizability to complex circuit designs.
@@
Although Wang et al. attempted to improve GNN generalizability by adopting the notion of graph contrastive learning [17], it involves nontrivial computational costs and thus worsens the GNN scalability. In this work, we aim to devise a graph learning model that is both scalable and generalizable for circuit designs, with the aid of a novel gated self-attention module applied on hop-wise features per node.
III The Proposed Approach
@@
Figure 2: (a) An overview of the two major phases of HOGA; (b) Our introduced gated self-attention module for hop-wise feature aggregation.
@@
Problem formulation. Given a circuit graph, we follow the common GNN settings that there are two inputs: graph adjacency matrix $A$ and node feature matrix $X$. Our goal is to build a model $\mathcal{M}$ for learning high-quality node representations $Y=\mathcal{M}(A,X)$, which can be utilized for a broad spectrum of circuit problems.
@@
Figure 2(a) gives an overview of our proposed approach, HOGA, which consists of two major phases. During the first phase, HOGA computes hop-wise features by iteratively multiplying the adjacency matrix with the node feature matrix, as described in Section III-A. Note that this step can be finished in advance, allowing HOGA training with complexity independent of the graph structure. The second phase trains a novel gated self-attention module on hop-wise features to learn node representations. This module enables HOGA to adaptively learn critical and high-order features from different hops per node, as further illustrated in Section III-B. Consequently, HOGA is adaptive to different structures across various circuit designs while simultaneously benefiting from high parallelism for distributed training.
III-A Hop-Wise Feature Generation
@@
Learning from neighborhood information is crucial for GNNs to capture critical graph structures. As depicted in Figure 1(b), traditional message-passing GNNs recursively aggregate features from neighbors, which leads to their poor scalability on large graphs. In contrast, we adopt a coarse-grained message-passing scheme based on hop-wise feature aggregation. To this ends, our first step is to generate hop-wise features.
@@
Given a graph $\mathcal{G}$ with its adjacency matrix $A\in\mathbb{R}^{n\times n}$ and node feature matrix $X\in\mathbb{R}^{n\times d}$, where $n$ and $d$ represent the number of nodes and the feature dimension respectively, we first normalize the adjacency matrix: $\hat{A}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, where $D$ is the node degree matrix. Next, we iteratively compute Equation (3) to generate hop-wise features for all nodes in parallel:
$X^{(k)}=\hat{A}X^{(k-1)},\;k=1,2,...,K$ (3)
where $X^{(0)}=X$, the $i$-th row of $X^{(k)}$ denotes the $k$-hop feature for node $i$, and $K$ is a hyperparameter to control the maximum number of hops for consideration.
@@
Consequently, for each $i\in\{1,2,...,n\}$, $\mathcal{X}_{i}$ comprises up to $K$-hop features of node $i$, which are then independently used to learn the corresponding node representation $Y_{i}$. Hence, there are no dependencies between different nodes for learning their representations, making it easy to scale HOGA through distributed training. It is noteworthy that the time required for generating hop-wise features is generally negligible compared to the overall training time, as empirically confirmed in Section IV-B.
@@
While there are a few prior arts (e.g. SIGN [7]) augmenting node features by adopting a similar approach to Equation (3), they simply train a multi-layer perceptron (MLP) model on augmented features. In contrast, we consider learning node representations through a hop-wise feature aggregation scheme, which is built upon a novel gated self-attention module introduced in the following section.
III-B Hop-Wise Aggregation with Gated Self-Attention
Since HOGA takes $\mathcal{X}_{i}\in\mathbb{R}^{(K+1)\times d}$ as input and produces a representation $Y_{i}\in\mathbb{R}^{d}$ for every node $i$ independently, we omit the node index $(i)$ in the ensuing discussion, by simply denoting $\mathcal{X}_{i}$ as $H$ and $Y_{i}$ as $y$ for clarity.
@@ III-B Hop-Wise Aggregation with Gated Self-Attention
A straightforward way of producing $y$ is to accumulate hop-wise features in $H$, i.e., $y=\sum_{k=0}^{K}H_{k}$. However, this approach has two fundamental limitations: (1) it fails to capture high-order feature interactions among different hop neighbors, resulting in its incapability of learning high-order structures that are critical for many EDA problems (e.g., functional reasoning), as discussed in Section II-B. (2) it uniformly combines features from different hops and thus cannot identify and focus on important hop-wise features per node.
To address the aforementioned limitations, let us first consider a simple gated layer as follows:
$U=HW_{U},V=HW_{V}$ (5)
$\hat{H}=U\odot V$ (6)
@@
where $W_{U},W_{V}\in\mathbb{R}^{d\times d}$ are trainable weight matrices, and $\odot$ denotes the element-wise product. By substituting Equation (5) into Equation (6), we can derive that $\hat{H}_{k}$ captures the second-order interaction $(H_{k}W_{U})\odot(H_{k}W_{V})$ for any hop $k\in\{0,1,...,K\}$. However, this also indicates that Equation (6) fails to capture interactions among different hop-wise features, i.e., $(H_{k}W_{U})\odot(H_{j}W_{V})$ with $k\neq j$. To tackle this limitation, we introduce a gated self-attention layer as follows:
$S=softmax(QK^{T})$ (7)
$\hat{H}=U\odot(SV)$ (8)
@@
where $Q=HW_{Q},K=HW_{K}$, $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ are trainable weight matrices, and $S$ is the self-attention matrix widely used in Transformer-based models [16]. Through substituting Equation (5) into Equation (8), we can derive that $\hat{H}_{k}=(H_{k}W_{U})\odot(\sum_{j=0}^{K}S_{k,j}H_{j}W_{V})=\sum_{j=0}^{K}S_{k,j}\;(H_{k}W_{U})\odot(H_{j}W_{V})$, which captures second-order interactions on different hop-wise features. By stacking more layers, the output $\hat{H}$ naturally captures higher-order feature interactions from different hops, which correspond to higher-order structures on the input graph $\mathcal{G}$.
@@
Figure 3: (a) An example of And-Inverter Graph, where each node represents a primary input/output (PI/PO) or an AND gate. The dashed arrows denote inverted edges; (b) An overview of QoR prediction. Our approach only replaces the GCN model in OpenABC-D with HOGA; (c) An overview of functional reasoning. Our approach solely substitutes HOGA for the GraphSAGE model in Gamora.
Implementation details. To improve the training stability of HOGA, we add LayerNorm and ReLU to Equation (8) in our implementation, resulting in the following:
$\hat{H}=\text{ReLU}(\text{LayerNorm}(U\odot(SV)))$ (9)
@@
After obtaining $\hat{H}=[\hat{H}_{0},\hat{H}_{1},...,\hat{H}_{K}]^{T}\in\mathbb{R}^{(K+1)\times d}$, we exploit an attentive readout function as per Equation (10) to produce the final node representation $y$:
$\begin{split}c_{k}&=\frac{exp(\alpha^{T}(\hat{H}_{0}||\hat{H}_{k}))}{\sum_{j=1}^{K}exp(\alpha^{T}(\hat{H}_{0}||\hat{H}_{j}))}\\ y&=\hat{H}_{0}+\sum_{k=1}^{K}c_{k}\hat{H}_{k}\end{split}$ (10)
@@
where $\alpha\in\mathbb{R}^{2d}$ is a trainable vector, ||||| | denotes the concatenation operator, and $c_{k}$ represents an attention score to measure the importance of the $k$-hop feature $\hat{H}_{k}$ to the final node representation $y$. In this way, HOGA can identify and adaptively aggregate critical features from different hops to produce high-quality node representations, which are then used for downstream tasks.
III-C Complexity analysis of HOGA

### 10 Apr 2024 v4

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 61st ACM/IEEE Design Automation Conference; June 23–27, 2024; San Francisco, CA, USAbooktitle: 61st ACM/IEEE Design Automation Conference (DAC ’24), June 23–27, 2024, San Francisco, CA, USAdoi: 10.1145/3649329.3657386isbn: 979-8-4007-0601-1/24/06
1. Introduction
@@
Recent years have seen a surge of interest in machine learning (ML) for electronic design automation (EDA), which holds great potential in achieving faster design closure and minimizing the need for extensive human supervision (huang-mleda-survey-todaes2021, ). In particular, graph neural networks (GNNs) have become increasingly popular in the EDA community due to their ability to encode graph-structured data such as gate-level netlists into compact representations, which can be used for a multitude of downstream EDA applications, including quality of results (QoR) prediction and functional reasoning (sanchez2023comprehensive, ; wu2023gamora, ).
@@
However, scaling GNN training to large graphs is a notoriously challenging problem, which poses a serious concern on the practical benefit of GNNs on large-scale EDA problems. On the one hand, unlike common datasets on social networks and molecular graphs, which consist of either a few large graphs or a large number of small graphs, the circuit datasets may contain numerous large graphs. For instance, the OpenABC-D benchmark provides $870$k gate-level netlists, where each netlist consists of up to $240$k logic gates (chowdhury2021openabc, ). Thus, training GNNs on such a large-scale circuit dataset is even more challenging than other graph-based applications. On the other hand, modern GNN models are built upon a message-passing paradigm, which learns representations through a recursive node-wise aggregation scheme shown in Figure 1(b). As a consequence, it is nontrivial to perform efficient distributed GNN training due to the node dependencies in a graph structure.
@@
Apart from the scalability challenge, it is also underexplored how to make GNNs generalizable across different circuit designs. Although there are many customized GNNs previously proposed for various EDA applications, their model backbones mainly follow classic GNNs such as GCN (kipf2016semi, ) and GraphSAGE (hamilton2017inductive, ), which are not necessarily suitable for circuit problems. Consider a task of identifying functional blocks within circuits (wu2023gamora, ). As distinct functional blocks may have different depths, the number of hops to be considered varies across nodes, which cannot be easily captured by common GNNs. Moreover, the high-order structures of functional blocks are also important yet ignored by the aforementioned GNN models. As a result, existing GNNs for EDA tasks often struggle to learn the intrinsic and critical information from complex circuit graphs, resulting in limited generalizability to unseen designs.
@@
Figure 1. Comparison of HOGA and prior GNNs — (a) An example graph for illustration; (b) GNN computation graph; (c) Computation graph of our proposed approach, HOGA.
@@
In literature, improving the scalability and enhancing the generalizability of GNNs on circuits are largely viewed as two orthogonal directions and seldomly explored together. Previous GNNs either incur massive communication overhead among GPUs, or rely on heuristic graph sampling algorithms that may lose critical structural information (shao2022distributed, ), which in turn degrades the model generalizability. In the context of learning generalizable GNNs for circuit problems, Wang et al. (wang2022functionality, ) leveraged graph contrastive learning to pretrain GNNs and adopted multiple neighbor aggregation functions to learn the functionalities of logic gates, which comes with additional computational costs and thus worsens the model scalability.
@@
Motivated by the limitations of conventional GNNs on circuit designs, we propose a novel hop-wise attention approach, named HOGA, to improve both the scalability and generalizability of circuit representation learning. As shown in Figure 1(c), HOGA adopts a hop-wise aggregation scheme, which precomputes the hop-wise features and only uses those features to learn node representations through a gated self-attention module. Since the final representation per node depends solely on its own hop-wise features, there are no dependencies between different nodes during training, making it easy to scale HOGA training in a distributed manner. Moreover, the gated self-attention module enables HOGA to adaptively capture critical high-order structures from different hops per node, rendering it generalizable across different circuit designs.
@@
Notably, HOGA is a flexible approach for learning circuit representations, which can be integrated with other customization techniques previously proposed for various downstream EDA tasks. To demonstrate the viability and flexibility of HOGA, we consider two representative circuit problems: (1) QoR prediction – We focus on predicting the optimized gate count after logic synthesis on the OpenABC-D benchmark, which is generated from various circuit designs as well as synthesis recipes, and is one of the largest open-sourced circuit datasets; (2) Functional reasoning – We follow the most challenging setting in Gamora (wu2023gamora, ) by identifying functional blocks on gate-level netlists after technology mapping. Our experiments show that HOGA not only outperforms prior GNNs on unseen designs, but is also efficient for distributed training. We summarize our main technical contributions as follows:
@@
$\bullet$ To the best of our knowledge, we are the first to introduce a scalable and generalizable model for circuit representation learning, which is achieved by a novel hop-wise graph attention scheme.
@@
$\bullet$ By precomputing hop-wise features, HOGA avoids recursive neighbor aggregation, which enables HOGA to learn each node representation independently and facilitates massive parallelization for distributed training. As a result, the training time of HOGA almost linearly decreases with an increase in computing resources.
@@
2. Preliminary and Motivation
2.1. Graph Neural Networks at Scale
@@
(1) $m_{i}^{(l)}=f^{(l)}(\{x_{j}^{(l-1)}\;|\;j\in\mathcal{N}(i)\}),\;x_{i}^{(l)}=g^{(l)}(m_{i}^{(l)},x_{i}^{(l-1)})\vspace{-5pt}$
@@
where $x_{i}^{(l)}$ denotes the node feature vector at the $l$-th layer, $\mathcal{N}(i)$ represents the set of neighbors of node $i$, $f$ can be any permutation-invariant function (e.g., mean-pooling), and the goal of function $g$ is to update node representations based on the aggregated features from neighbors. After stacking $L$ GNN layers, the output features $\{x_{i}^{(L)}\;|\;i\in\mathcal{V}\}$ serve as the final node representations. However, let $d$ be the average node degree, there are $O(d^{L})$ nodes required to obtain $x_{i}^{(L)}$, which increases exponentially with the number of layers and is known as the “neighbor explosion” issue. Besides, the graph dependencies also incur significant communication overhead and work imbalance, hindering efficient distributed GNN training. To tackle both challenges, numerous efforts have been devoted to improve GNN training efficiency by adopting different sampling strategies (hamilton2017inductive, ; chen2018fastgcn, ; zeng2019graphsaint, ). While these methods demonstrate promising results on social networks, we argue that they are unsuitable for circuits as the sampling algorithms may entirely break the design functionality and lead to poor accuracy. This is empirically confirmed in Section 4.3. In contrast, we introduce a hop-wise aggregation scheme that independently learns each node representation based on its own hop-wise features, which renders our model embarrassingly parallel and greatly facilitates distributed training.
2.2. Generalizable Graph Learning in EDA
@@
Since realistic circuit designs may originate from distinct domains, generalization capability is crucial for deploying graph learning models on circuits. To this end, Ustun et al. proposed a customized GNN model by distinguishing predecessors and successors in a graph, which demonstrates promising generalizability on learning operation mapping patterns (ustun2020dsage, ). Later, Zhang et al. (zhang2020grannite, ) and Guo et al. (guo2022timing, ) introduced customized GNNs tailored for power inference and timing prediction tasks respectively, via sequentially updating node representations. More recently, Wu et al. presented a multi-task graph learning framework for functional reasoning. While these methods showcase promising results on the respective tasks, their GNN backbones are built upon the conventional message-passing paradigm, which cannot capture critical high-order structures formed by multiple nodes. Notably, many functional blocks (e.g., full adders) are essentially high-order structures, which are crucial for many circuit problems such as functional reasoning (wu2023gamora, ). Therefore, the aforementioned methods still cannot capture the intrinsic circuit information, leading to either restrictions to specific tasks or limited generalizability to complex circuit designs.
@@
Although Wang et al. attempted to improve GNN generalizability by adopting the notion of graph contrastive learning (wang2022functionality, ), it involves nontrivial computational costs and thus worsens the GNN scalability. In this work, we aim to devise a graph learning model that is both scalable and generalizable for circuit designs by using a gated self-attention module on hop-wise features per node.
3. The Proposed Approach
@@
Figure 2. (a) An overview of the two major phases of HOGA; (b) Our introduced gated self-attention module.
@@
Problem formulation. Given a graph adjacency matrix $A$ and a node feature matrix $X$, our goal is to build a model $\mathcal{M}$ for learning high-quality node representations $Y=\mathcal{M}(A,X)$, which can be utilized for a broad spectrum of circuit problems.
@@
Figure 2(a) gives an overview of HOGA that consists of two major phases. During the first phase, HOGA computes hop-wise features by iteratively multiplying the adjacency matrix with the node feature matrix, as described in Section 3.1. Note that this step can be finished in advance, allowing HOGA training with complexity independent of the graph structure. The second phase trains a gated self-attention module to capture high-order interactions among hop-wise features, as illustrated in Section 3.2. Consequently, HOGA is generalizable to different circuit designs while simultaneously benefiting from high parallelism for distributed training.
3.1. Hop-Wise Feature Generation
@@
As depicted in Figure 1(b), traditional message-passing GNNs recursively aggregate features from neighbors, which leads to their poor scalability on large graphs. In contrast, we adopt a coarse-grained message-passing scheme based on hop-wise feature aggregation. To this ends, our first step is to generate hop-wise features.
@@
Given the adjacency matrix $A\in\mathbb{R}^{n\times n}$ and node feature matrix $X\in\mathbb{R}^{n\times d}$, where $n$ and $d$ represent the number of nodes and the feature dimension respectively, we first normalize the adjacency matrix: $\hat{A}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, where $D$ is the node degree matrix. Next, we generate hop-wise features by iteratively computing Equation (2), where $X^{(0)}=X$ and $K$ denotes the number of hops.
(2) $X^{(k)}=\hat{A}X^{(k-1)},\;k=1,2,...,K$
@@
Consequently, for each $i\in\{1,2,...,n\}$, $\mathcal{X}_{i}$ comprises up to $K$-hop features of node $i$, which are then independently used to learn the corresponding node representation $Y_{i}$. Hence, there are no dependencies between different nodes, making it easy to scale HOGA through distributed training. It is noteworthy that the time required for generating hop-wise features is generally negligible compared to the overall training time, as empirically confirmed in Section 4.2.
@@
While there are a few prior arts (e.g. SIGN (frasca2020sign, )) augmenting node features by adopting a similar approach to Equation (2), they simply train a multi-layer perceptron (MLP) model on augmented features. In contrast, we consider learning node representations through a hop-wise feature aggregation scheme, which is built upon a novel gated self-attention module introduced in the following section.
3.2. Hop-Wise Gated Attention
@@
Since HOGA takes $\mathcal{X}_{i}\in\mathbb{R}^{(K+1)\times d}$ as input and produces a representation $Y_{i}\in\mathbb{R}^{d}$ for every node $i$ independently, we omit the node index $(i)$ in the ensuing discussion, by simply denoting $\mathcal{X}_{i}$ as $H$ and $Y_{i}$ as $y$ for clarity. A straightforward way of producing $y$ is to accumulate hop-wise features in $H$, i.e., $y=\sum_{k=0}^{K}H_{k}$. However, this approach has two flaws: (1) it fails to capture high-order feature interactions among different hop neighbors, resulting in its limited expressivity of learning high-order circuit structures. (2) it uniformly combines features from different hops and thus cannot identify and focus on important hop-wise features per node. To address those limitations, let us first consider a simple gated layer:
(4) $U=HW_{U},V=HW_{V},\;\hat{H}=U\odot V\vspace{-3pt}$
@@
where $W_{U},W_{V}\in\mathbb{R}^{d\times d}$ are trainable weight matrices, and $\odot$ denotes the element-wise product. We can derive from Equation (4) that $\hat{H}_{k}$ captures the second-order interaction $(H_{k}W_{U})\odot(H_{k}W_{V})$ for every hop $k\in\{0,1,...,K\}$. However, this also means Equation (4) fails to capture interactions among different hop-wise features, i.e., $(H_{k}W_{U})\odot(H_{j}W_{V})$ with $k\neq j$. To tackle this issue, we introduce a gated self-attention layer in the following:
(5) $S=softmax(QK^{T}),\;\hat{H}=U\odot(SV)\vspace{-3pt}$
@@
where $Q=HW_{Q},K=HW_{K}$, $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ are trainable weight matrices, and $S$ is the self-attention matrix widely used in Transformer (vaswani2017attention, ). Based on Equations (4) and (5), we can derive that $\hat{H}_{k}=(H_{k}W_{U})\odot(\sum_{j=0}^{K}S_{k,j}H_{j}W_{V})=\sum_{j=0}^{K}S_{k,j}\;(H_{k}W_{U})\odot(H_{j}W_{V})$, which captures second-order interactions on different hop-wise features. By stacking more layers, the output $\hat{H}$ naturally captures higher-order feature interactions from different hops, which correspond to higher-order structures on the input circuit graph.
@@
Figure 3. (a) An example of And-Inverter Graph, where each node represents a primary input/output (PI/PO) or an AND gate. The dashed arrows denote inverted edges; (b) An overview of QoR prediction in OpenABC-D; (c) An overview of functional reasoning in Gamora — We solely replace GNNs with HOGA for learning circuit representations in both tasks.
@@
Implementation details. To improve the training stability of HOGA, we add LayerNorm and ReLU to Equation (5) in our implementation, i.e., $\hat{H}=\text{ReLU}(\text{LayerNorm}(U\odot(SV)))$. After obtaining $\hat{H}=[\hat{H}_{0},\hat{H}_{1},...,\hat{H}_{K}]^{T}\in\mathbb{R}^{(K+1)\times d}$, we adopt an attentive readout scheme to produce the final node representation $y$:
(6) $\begin{split}c_{k}=\frac{exp(\alpha^{T}(\hat{H}_{0}||\hat{H}_{k}))}{\sum_{j=1}^{K}exp(\alpha^{T}(\hat{H}_{0}||\hat{H}_{j}))},\;y=\hat{H}_{0}+\sum_{k=1}^{K}c_{k}\hat{H}_{k}\end{split}$
@@
where $\alpha\in\mathbb{R}^{2d}$ is a trainable vector, ||||| | denotes the concatenation operator, and $c_{k}$ represents an attention score to measure the importance of the $k$-hop feature $\hat{H}_{k}$ to the final node representation $y$. In this way, HOGA can identify and adaptively aggregate critical features from different hops to produce high-quality node representations, which are then used for downstream circuit tasks.
3.3. Complexity analysis of HOGA